OpenStudy (anonymous):

. A person travels by car from one city to another with different constant speeds between pairs of cities. She drives for 30.0 min at 80.0 km/h, 12.0 min at 100 km/h, and 45.0 min at 40.0 km/h and spends 15.0 min eating lunch and buying gas. (a) Determine the average speed for the trip. (b) Determine the distance between the initial and final cities along the route.

OpenStudy (radar):

Distance is the sum of: 40 km (.5*80), 20 (.2*100), and 30 (.75*40). Distance equals 40 + 20 + 30 = 90 km Total distance = 90 km. Total time (including lunch break) is 1/2 hr + 1/5 hr +3/4 hr or .5 + .2 + .75 = 1.45 hrs Average speed for trip: 90/1.45 =62.07 km/kr

OpenStudy (radar):

*hr